OpenStudy (anonymous):

(I need help, this is three parts :p) While Playing Yahtzee and rolling 5 dice, Mei gets the result shown at right (the numbers are 2, 4, 4, 4, and 1). Mei decides to keep the three fours and reroll the other 2 dice. a. what is the probability that Mei will have 5 of a kind? b. What is the probability that she will have 4 of a kind? c. what is the probability that she will have exactly three 4's?

OpenStudy (anonymous):

The easiest way to do this question is to ignore the dice that have already been rolled. And instead consider them as just two dice. a. what is the probability Mei will roll 2 fours b. what is the probability Mei will roll 1 four c. what is the probability Mei will roll no fours

OpenStudy (anonymous):

Would you be able to work that out?

OpenStudy (anonymous):

a. 1/6 times 1/6 times b. 1/6 c. 0

OpenStudy (anonymous):

a. 1/6 times 1/6

OpenStudy (anonymous):

I think maybe...@riccyscaduto a. 1/6 b. c. 5/6

OpenStudy (anonymous):

That's close, but you need to consider 2 dice. So, the probability of rolling a 4 on a single dice would be 1/6 When there are two or more occurrences you need to multiply the probabilities

OpenStudy (anonymous):

1/6 chance of rolling a four on the first and 1/6 chance of rolling a four on the second would be?

OpenStudy (anonymous):

yeah but if we figure out rolling a four on 2 isn't it 2/12? that simplifies to 1/6....

OpenStudy (anonymous):

a. 1/36 b. 1/6 c. 0

OpenStudy (anonymous):

You add probabilities when there are more than one way of reaching the same outcome. For this you need to multiply\[\frac{ 1 }{ 6 } * \frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

oh yeah so that would be 1/36

OpenStudy (anonymous):

Yeah, that's right. b. is marginally more complicated. There are two ways of rolling 1 four. rolling a 1/2/3/5/6 with the first dice and rolling a 4 on the second or rolling a 4 on the first dice and rolling 1/2/3/5/6 on the second

OpenStudy (anonymous):

so \[\frac{ 5 }{ 6 } * \frac{ 1 }{ 6 } \] and\[\frac{ 1 }{ 6 } * \frac{ 5 }{ 6 } \]

OpenStudy (anonymous):

Does that make sense? I know it's a little messy

OpenStudy (anonymous):

wait so its 5/36?

OpenStudy (anonymous):

for one of them, yes and this is what I mentioned earlier, when there are two ways of getting the same outcome you and them so \[\frac{ 5 }{ 36 } + \frac{ 5 }{ 36 }\]

OpenStudy (anonymous):

you add them*

OpenStudy (anonymous):

so it would be 10/36 or 5/18?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

C is similar to a except it's the chance of not rolling a 4

OpenStudy (anonymous):

and the answer to c is the same though

OpenStudy (anonymous):

the chance of not rolling a 4 on each dice is 5/6 so the answer would be \[\frac{ 5 }{ 6 } * \frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

so 25/36 and it can't be reduced.

OpenStudy (anonymous):

yes, so that's all 3 answers

OpenStudy (anonymous):

thank you so much!